# A Catalog of $\sim 146,000$ Radio Galaxies up to $z\sim 5$

### v1

@@ Abstract
We present the results of optical identifications for Very Large Array Sky Survey (VLASS) radio sources employing the Ultraviolet Near Infrared Optical Northern Survey (UNIONS). A cross-match between UNIONS and VLASS Epoch 2 catalogs yields $146,212$ radio galaxies down to $r=24.5$ mag over a wide area of $\sim 4,200\mathrm{\ deg^{2}}$ in the northern hemisphere. We perform $g-$dropout selections and find $\gtrsim 200$ sources at $z\sim 4$ optimistically. Of 63,019 sources with valid photometric redshifts, 8,692 are at ${z}_{\mathrm{photo}}\geq 1$ and 1,171 are at ${z}_{\mathrm{photo}}\geq 2$. Based on spectral luminosities at 1.4 and 3 GHz using the valid UNIONS photo-z, we identify $\sim 49,000$ radio-loud AGNs (RLAGNs) with ${L}_{\mathrm{1.4GHz}}>10^{24}\mathrm{\ W\ Hz^{-1}}$, and all radio galaxies at $z\geq 1$ are RL. Adopting radio loudness instead, 138,266 out of 146,212 UNIONS-VLASS radio galaxies are RL. By virtue of the wide area and medium depth provided by UNIONS, our catalog greatly increases the number counts of RLAGNs at $z>1$. We further cross-match the UNIONS-VLASS catalog with LOw-Frequency ARray Two-metre Sky Survey (LoTSS) at 144 MHz and Faint Images of the Radio Sky at Twenty-cm (FIRST) at 1.4 GHz, yielding 101,671 UNIONS-VLASS-LoTSS, 79,638 -FIRST, and 64,672 -LoTSS-FIRST sources, respectively. This multifrequency radio dataset reveals sources of various spectral shapes, including the steep spectrum of aged populations, the peaked spectrum of young populations, and the upturned spectrum that might be associated with transient sources. The UNIONS-VLASS radio galaxy candidates will be covered by the Euclid wide survey, bringing about legacy values to benefit multi-faceted studies related to radio galaxies and beyond.
@@
facilities: VLA, CFHT, Pan-STARRS, Subaru HSC, and LOFARsoftware: astropy (Astropy Collaboration et al., 2013, 2018); astroquery (Ginsburg et al., 2019)
{CJK*}
UTF8gkai
@@ 1 Introduction
Supermassive black holes (SMBHs) are believed to ubiquitously reside in the center of most galaxies (Kormendy & Ho, 2013; McConnell & Ma, 2013), and a tight correlation between the mass of a SMBH and the properties of its host galaxy has been observed since the 1990s (Magorrian et al., 1998; Ferrarese & Merritt, 2000; Gebhardt et al., 2000; Tremaine et al., 2002). The tight correlation implies that SMBHs grow in tandem with the parameters of their host galaxies, so-called “co-evolution”. How SMBHs co-evolve with their host galaxies and how they impact their host galaxy evolution are long-standing questions in modern astronomy.
@@
One key phenomenon to shape such co-evolution is called AGN feedback. Such a feedback can lead to the outflowing and/or heating of the cold gas, suppressing the host galaxy star formation and self-regulating the growth of the SMBH (Fabian, 2012; Alexander & Hickox, 2012; Bischetti et al., 2022; Harrison & Ramos Almeida, 2024; Talbot et al., 2024; Huško et al., 2025), although a compression of cold gas that enhances the star formation has also been observed (Shin et al., 2019; Duncan et al., 2023). One mechanism of AGN feedback is through the coupling between the host galaxy ISM and the relativistic plasma, which is known as radio jets, launched from the central SMBH. This radio/jet-mode feedback is considered the primary mode of AGN feedback in radio-loud active galactic nuclei (RLAGNs).
@@
RLAGNs are often selected via the observed radio loudness $\mathcal{R}_{\mathrm{obs}}(={F}_{\mathrm{radio}}/{F}_{\mathrm{optical}}>10)$ defined as the ratio of radio-to-optical flux densities (Kellermann et al., 1989; Balokovic et al., 2012), or via the 1.4 GHz spectral luminosity ${L}_{\mathrm{1.4GHz}}\gtrsim 10^{24}\mathrm{\ W\ Hz^{-1}}$; Peacock et al. 1986; Condon 1989; Tadhunter 2016; Padovani et al. 2017). In the local Universe, RLAGNs are found to preferentially reside in early-type galaxies correlated with high stellar masses, low star formation rates, and radiatively inefficient AGNs (Best et al. 2005; Heckman & Best 2014; Gürkan et al. 2015; Pierce et al. 2022; however, see also Ramos Almeida et al. 2012; Pierce et al. 2023). At higher redshifts ($z>1$), there is an increasing number of RLAGNs found to be radiatively efficient (Eddington ratio ${\lambda}_{\mathrm{Edd}}\gtrsim 3\times 10^{-2}$; Merloni & Heinz, 2008; Ichikawa et al., 2021) and their host galaxies can be star-forming or even starburst galaxies (Drouart et al., 2014; Zhu et al., 2023; Zhong et al., 2024). These high-redshift radio galaxies (HzRGs) at $z>1$ are found tightly correlated with galaxy mergers (Chiaberge et al., 2015; Zhong et al., 2023; Heckman et al., 2024), suggestive of merger-triggered star formation and launching of radio jets. Therefore, RLAGNs across cosmic time, especially HzRGs, are valuable laboratories to study the interplay between the host galaxy evolution, SMBH growth, and radio jets, as well as the jet-accretion disc connections.
@@
The rareness of RLAGNs ($10-20\%$ among all AGN populations; Kellermann et al. 1989; Urry & Padovani 1995; Ivezić et al. 2002; Chiaberge & Marconi 2011) requires large surveys to increase the sample size to statistically understand RLAGNs, their host galaxies, and the connections in between. Over the past decades, following the groundbreaking Cambridge Catalogue of Radio Sources (e.g., Kellermann et al., 1969), many radio sky surveys have been conducted centered around $\sim 1$ GHz, including the National Radio Astronomy Observatory (NRAO) Very Large Array (VLA) Sky Survey (NVSS; 1.4 GHz; Condon et al., 1998) and VLA Faint Images of the Radio Sky at Twenty-cm (FIRST; 1.4 GHz; Becker et al., 1995), and Australian Square Kilometre Array Pathfinder (ASKAP) Evolutionary Map of the Universe (EMU; 944 MHz; Norris et al., 2011) and Rapid ASKAP Continuum Survey (RACS; $\sim 900-1600$ MHz; McConnell et al., 2020). The vast sky coverages ($\mathrm{>10,000\ deg^{2}}$) of these surveys offer great opportunities to build up an unbiased sample to elevate the statistical importance of studies related to RLAGNs.
@@
Wide-area optical surveys are necessary to identify the optical counterparts of radio sources and determine their redshifts, either photometrically or spectroscopically. Using the Sloan Digital Sky Survey (SDSS; York et al. 2000), Ivezić et al. (2002) and Helfand et al. (2015) identified the optical counterparts for the FIRST radio core sources at 1.4 GHz, adopting a positional cross-match between the two catalogs. Due to the shallow optical sensitivity of SDSS $r$-band imaging ($r<22.2$ mag), the fraction of matches (${f}_{\mathrm{matches}}$) is only $\sim 30\%$, and a significant fraction of optically faint host galaxies may have been missed. Using the Subaru Hyper Suprime-Cam (HSC) Subaru Strategic Program (SSP; Aihara et al., 2018), Yamashita et al. (2018) reaches ${f}_{\mathrm{matches}}=50.6\%$ optimistically for FIRST sources in the HSC field thanks to the depth of HSC reaching a limiting magnitude of 25.9 in i-band. Despite an improvement in associating radio sources with faint objects, elevating the fairness of the radio galaxy catalog, the survey area is $\mathrm{1,400\ deg^{2}}$ such that the catalog size is limited.
@@
Other than the 1.4 GHz continuum, radio galaxy catalogs have been constructed centering at 3 GHz and 144 MHz. Based on the largest sample of uniformly selected and spectrally confirmed quasars provided by SDSS, Arnaudova et al. (2024) identified 3,697 RL and 111,132 RQ sources in LOw-Frequency ARray (LOFAR) observations at 144 MHz via the radio loudness dichotomy. With 384 h observations of the upgraded Karl G. Jansky VLA, the VLA-COSMOS 3 GHz Large Project (Smolčić et al., 2017) identifies 10,830 radio sources in the COSMOS field. This project is equipped with a sensitivity better than that of FIRST by almost two orders of magnitude, greatly exploring the RQ populations. These catalogs are advantageous explorations in the faint end of radio galaxies; however, they do not significantly expand the number counts of RLAGNs.
@@
To improve both the sample size of RLAGNs and the unbiasedness of host galaxy populations, we search for UNIONS (Gwyn et al., 2025) optical counterparts for VLASS (Lacy et al., 2020) radio sources at 3 GHz. UNIONS reaches a depth of 24.7 mag in r-band and covers an extremely wide area of $\sim 5,000\mathrm{\ deg^{2}}$ that fully overlaps with VLASS, bridging the gap between the shallow imaging of SDSS and the deep but slightly limited sky coverage of HSC-SSP. VLASS is the state-of-the-art radio survey operated at 3 GHz in terms of spatial resolution, sensitivity, and total survey area. The cross-match between these two catalogs will produce one of the largest RLAGN catalogs covering a wide redshift range. Furthermore, both catalogs are ongoing projects, ensuring a great scalability of the UNIONS-VLASS catalog.
@@
In this paper, we use radio galaxies for sources cross-matched between the UNIONS optical and VLASS radio catalogs, and define HzRGs as radio galaxies at $z>1$. This paper is organized as follows. In Section 2, we introduce the optical and radio source catalogs used to build the radio galaxy catalog. The cross-matches between the catalogs and the statistical results of our radio galaxy catalog are presented in Section 3. We examined the optical and radio number counts of UNIONS-VLASS radio galaxies in Section 4, and we further presented the selection of RLAGNs based on the observed radio loudness and spectral luminosity. We show radio galaxy classifications using multifrequency radio datasets and discuss the high-z sources based on the dropout technique in Section 5. A summary is presented in Section 6. Throughout this paper, we assume a $\mathrm{\Lambda CDM}$ cosmology with $\Omega_{m}=0.309$, $\Omega_{\Lambda}=0.691$, and $H_{0}=\mathrm{67.7\ km\ s^{-1}\ Mpc^{-1}}$ (Planck Collaboration et al., 2016).
@@ 2.1 UNIONS
An elevation of the statistical importance of our radio galaxies catalog requires a wide ($>1000$ deg2) and deep ($r>22$ mag) optical survey. The Ultraviolet Near-Infrared Optical Northern Survey (UNIONS; Gwyn et al. 2025), as a science collaboration between wide imaging surveys that cover the northern hemisphere, provides an excellent opportunity to abundantly discover the optical counterparts of radio sources. The UNIONS integrates multiband (five broadbands, $u$, $g$, $r$, $i$, $z$) observations from Canada-France-Hawaii Telescope/Canada-France Imaging Survey (CFHT/CFIS, $u+r$), Subaru/Waterloo-Hawaii-IfA G-band Survey (WHIGS, g), The Panoramic Survey Telescope & Rapid Response System (Pan-STARRS, i), and Subaru/Wide Imaging with Subaru HSC of the Euclid Sky (WISHES, z). The main UNIONS footprint is located at a declination (DEC.) $>30^{\circ}$ and a galactic latitude of $|b|>25^{\circ}$, enclosing a total area of $4,861\mathrm{\ deg^{2}}$. Extended $u-$ and $z-$ (WISHES+) observations are being conducted to cover the sky area at $\mathrm{DEC.}<30^{\circ}$, elevating the sky coverage of UNIONS. Albeit UNIONS is motivated by the Euclid wide survey (Euclid Collaboration et al., 2022), efforts have been devoted to maintaining UNIONS as an independent survey with scientific values beyond what will be brought about by the Euclid.
@@
The $r-$band observation of CFHT/CFIS has an excellent median seeing of $0.69\arcsec$ and reaches a surface brightness limit of 28.4 $\mathrm{mag/arcsec^{2}}$ for extended sources. It is set as the reference catalog by which the elliptical aperture is defined for the forced photometry applied to all five UNIONS bands. Since the broadband observations of UNIONS are conducted under multiple telescopes, the Gaussian Aperture and Photometry ($\mathtt{GAaP}$; Kuijken et al. 2015) is adopted to produce Gaussianized images for raw imaging data collected from observations at different instruments. The forced photometry with Gaussian-weighted, elliptical apertures can then be properly used to perform homogeneous multiband photometric measurements simultaneously. The depth of UNIONS is characterized by the $10\sigma$ point source depth measured through a $2.0\arcsec$ aperture. The current median depth of UNIONS is $[u,g,r,i,{z}_{\mathrm{PS}},{z}_{\mathrm{HSC}}]=[23.7,24.5,24.2,23.8,23.3,23.3]$, to be compared with the original goal of $[u,g,r,i,z]=[23.6,24.5,23.9,23.6,23.4]$ (Gwyn et al., 2025).

### v2

@@ Abstract
We present the results of optical identifications for Very Large Array Sky Survey (VLASS) radio sources employing the Ultraviolet Near Infrared Optical Northern Survey (UNIONS). A cross-match between UNIONS and VLASS Epoch 2 catalogs yields $146,212$ radio galaxies down to $r=24.5$ mag over a wide area of $\sim 4,200\mathrm{\ deg^{2}}$ in the northern hemisphere. We perform $g$-dropout selections and find $\gtrsim 200$ sources at $z\sim 4$ optimistically. Of 63,019 sources with valid photometric redshifts, 8,692 are at ${z}_{\mathrm{photo}}\geq 1$ and 1,171 are at ${z}_{\mathrm{photo}}\geq 2$. Based on spectral luminosities at 1.4 and 3 GHz using the valid UNIONS photo-z, we identify $\sim 49,000$ radio-loud AGNs (RLAGNs) with ${L}_{\mathrm{1.4GHz}}>10^{24}\mathrm{\ W\ Hz^{-1}}$, and all radio galaxies at $z\geq 1$ are RL. Adopting radio loudness instead, 138,266 out of 146,212 UNIONS-VLASS radio galaxies are RL. By virtue of the wide area and medium depth provided by UNIONS, our catalog greatly increases the number counts of RLAGNs at $z>1$. We further cross-match the UNIONS-VLASS catalog with LOw-Frequency ARray Two-metre Sky Survey (LoTSS) at 144 MHz and VLA Faint Images of the Radio Sky at Twenty-cm (FIRST) at 1.4 GHz, yielding 101,671 UNIONS-VLASS-LoTSS, 79,638 -FIRST, and 64,672 -LoTSS-FIRST sources, respectively. This multifrequency radio dataset reveals sources of various spectral shapes, including the steep spectrum of aged populations, the peaked spectrum of young populations, and the upturned spectrum that might be associated with transient sources. The UNIONS-VLASS radio galaxy candidates will be covered by the Euclid wide survey, bringing about legacy values to benefit multi-faceted studies related to radio galaxies and beyond.
@@
facilities: VLA, CFHT, Pan-STARRS, Subaru HSC, and LOFARsoftware: astropy (astropy_collaboration_astropy_2013; astropy_collaboration_astropy_2018); astroquery (ginsburg_astroquery_2019)
@@ 1 Introduction
Supermassive black holes (SMBHs) are believed to ubiquitously reside in the center of most galaxies (kormendy_coevolution_2013; mcconnell_revisiting_2013), and a tight correlation between the mass of a SMBH and the properties of its host galaxy has been observed since the 1990s (magorrian_demography_1998; ferrarese_fundamental_2000; gebhardt_relationship_2000; tremaine_slope_2002). The tight correlation implies that SMBHs grow in tandem with the parameters of their host galaxies, so-called “co-evolution”. How SMBHs co-evolve with their host galaxies and how they impact their host galaxy evolution are long-standing questions in modern astronomy.
@@
One key phenomenon to shape such co-evolution is called AGN feedback. Such a feedback can lead to the outflowing and/or heating of the cold gas, suppressing the host galaxy star formation and self-regulating the growth of the SMBH (fabian_observational_2012; alexander_what_2012; bischetti_suppression_2022; harrison_observational_2024; talbot_simulations_2024; husko_effects_2025), although a compression of cold gas that enhances the star formation has also been observed (shin_positive_2019; duncan_jwsts_2023). One mechanism of AGN feedback is through the coupling between the host galaxy ISM and the relativistic plasma, which is known as radio jets, launched from the central SMBH. This radio/jet-mode feedback is considered the primary mode of AGN feedback in radio-loud active galactic nuclei (RLAGNs).
@@
RLAGNs are often selected via the observed radio loudness $\mathcal{R}_{\mathrm{obs}}(={F}_{\mathrm{radio}}/{F}_{\mathrm{optical}}>10)$ defined as the ratio of radio-to-optical flux densities (kellermann_vla_1989; balokovic_disclosing_2012), or via the 1.4 GHz spectral luminosity ${L}_{\mathrm{1.4GHz}}\gtrsim 10^{24}\mathrm{\ W\ Hz^{-1}}$; peacock_statistics_1986; condon_14_1989; tadhunter_radio_2016; padovani_active_2017). In the local Universe, RLAGNs are found to preferentially reside in early-type galaxies correlated with high stellar masses, low star formation rates, and radiatively inefficient AGNs (best_host_2005; heckman_coevolution_2014; gurkan_herschel-atlas_2015; pierce_agn_2022; however, see also ramos_almeida_are_2012; pierce_galaxy_2023). At higher redshifts ($z>1$), there is an increasing number of RLAGNs found to be radiatively efficient (Eddington ratio ${\lambda}_{\mathrm{Edd}}\gtrsim 3\times 10^{-2}$; merloni_synthesis_2008; ichikawa_wide_2021) and their host galaxies can be star-forming or even starburst galaxies (drouart_rapidly_2014; zhu_radio_2023; zhong_revisiting_2024). These high-redshift radio galaxies (HzRGs) at $z>1$ are found tightly correlated with galaxy mergers (chiaberge_radio_2015; zhong_revisiting_2023; heckman_mergers_2024), suggestive of merger-triggered star formation and launching of radio jets. Therefore, RLAGNs across cosmic time, especially HzRGs, are valuable laboratories to study the interplay between the host galaxy evolution, SMBH growth, and radio jets, as well as the jet-accretion disc connections.
@@
The rareness of RLAGNs ($10-20\%$ among all AGN populations; kellermann_vla_1989; urry_unified_1995; ivezic_optical_2002; chiaberge_origin_2011) requires large surveys to increase the sample size to statistically understand RLAGNs, their host galaxies, and the connections in between. Over the past decades, following the groundbreaking Cambridge Catalogue of Radio Sources (e.g., kellermann_spectra_1969), many radio sky surveys have been conducted centered around $\sim 1$ GHz, including the National Radio Astronomy Observatory (NRAO) Very Large Array (VLA) Sky Survey (NVSS; 1.4 GHz; condon_nrao_1998) and VLA Faint Images of the Radio Sky at Twenty-cm (FIRST; 1.4 GHz; becker_first_1995), and Australian Square Kilometre Array Pathfinder (ASKAP) Evolutionary Map of the Universe (EMU; 944 MHz; norris_emu_2011) and Rapid ASKAP Continuum Survey (RACS; $\sim 900-1600$ MHz; mcconnell_rapid_2020). The vast sky coverages ($\mathrm{>10,000\ deg^{2}}$) of these surveys offer great opportunities to build up an unbiased sample to elevate the statistical importance of studies related to RLAGNs.
@@
Wide-area optical surveys are necessary to identify the optical counterparts of radio sources and determine their redshifts, either photometrically or spectroscopically. Using the Sloan Digital Sky Survey (SDSS; york_sloan_2000), ivezic_optical_2002 and helfand_last_2015 identified the optical counterparts for the FIRST radio core sources at 1.4 GHz, adopting a positional cross-match between the two catalogs. Due to the shallow optical sensitivity of SDSS $r$-band imaging ($r<22.2$ mag), the fraction of matches (${f}_{\mathrm{matches}}$) is only $\sim 30\%$, and a significant fraction of optically faint host galaxies may have been missed. Using the Subaru Hyper Suprime-Cam (HSC) Subaru Strategic Program (SSP; aihara_hyper_2018), yamashita_wide_2018 reaches ${f}_{\mathrm{matches}}=50.6\%$ optimistically for FIRST sources in the HSC field thanks to the depth of HSC reaching a limiting magnitude of 25.9 in i-band. Despite an improvement in associating radio sources with faint objects, elevating the fairness of the radio galaxy catalog, the survey area is $\mathrm{1,400\ deg^{2}}$ such that the catalog size is limited.
@@
Other than the 1.4 GHz continuum, radio galaxy catalogs have been constructed centering at 3 GHz and 144 MHz. Based on the largest sample of uniformly selected and spectrally confirmed quasars provided by SDSS, arnaudova_exploring_2024 identified 3,697 RL and 111,132 RQ sources in LOw-Frequency ARray (LOFAR) observations at 144 MHz via the radio loudness dichotomy. With 384 h observations of the upgraded Karl G. Jansky VLA, the VLA-COSMOS 3 GHz Large Project (smolcic_vla-cosmos_2017) identifies 10,830 radio sources in the COSMOS field. This project is equipped with a sensitivity better than that of FIRST by almost two orders of magnitude, greatly exploring the RQ populations. These catalogs are advantageous explorations in the faint end of radio galaxies; however, they do not significantly expand the number counts of RLAGNs.
@@
To improve both the sample size of RLAGNs and the unbiasedness of host galaxy populations, we search for UNIONS (gwyn_unions_2025) optical counterparts for VLASS (lacy_karl_2020) radio sources at 3 GHz. UNIONS reaches a depth of 24.7 mag in r-band and covers an extremely wide area of $\sim 5,000\mathrm{\ deg^{2}}$ that fully overlaps with VLASS, bridging the gap between the shallow imaging of SDSS and the deep but slightly limited sky coverage of HSC-SSP. VLASS is the state-of-the-art radio survey operated at 3 GHz in terms of spatial resolution, sensitivity, and total survey area. The cross-match between these two catalogs will produce one of the largest RLAGN catalogs covering a wide redshift range. Furthermore, both catalogs are ongoing projects, ensuring a great scalability of the UNIONS-VLASS catalog.
@@
In this paper, we use radio galaxies for sources cross-matched between the UNIONS optical and VLASS radio catalogs, and define HzRGs as radio galaxies at $z>1$. This paper is organized as follows. In Section 2, we introduce the optical and radio source catalogs used to build the radio galaxy catalog. The cross-matches between the catalogs and the statistical results of our radio galaxy catalog are presented in Section 3. We examine the optical and radio number counts of UNIONS-VLASS radio galaxies in Section 4, and we further present the selection of RLAGNs based on the observed radio loudness and spectral luminosity. We show radio galaxy classifications using multifrequency radio datasets and discuss the high-z sources based on the dropout technique in Section 5. A summary is presented in Section 6. Throughout this paper, we assume a $\mathrm{\Lambda CDM}$ cosmology with $\Omega_{m}=0.309$, $\Omega_{\Lambda}=0.691$, and $H_{0}=\mathrm{67.7\ km\ s^{-1}\ Mpc^{-1}}$ (planck_collaboration_planck_2016).
@@ 2.1 UNIONS
An elevation of the statistical importance of our radio galaxies catalog requires a wide (>1000>1000 deg2) and deep ($r>22$ mag) optical survey. The Ultraviolet Near-Infrared Optical Northern Survey (UNIONS; gwyn_unions_2025), as a science collaboration between wide imaging surveys that cover the northern hemisphere, provides an excellent opportunity to abundantly discover the optical counterparts of radio sources. The UNIONS integrates multiband (five broadbands, $u$, $g$, $r$, $i$, $z$) observations from Canada-France-Hawaii Telescope/Canada-France Imaging Survey (CFHT/CFIS, $u+r$), Subaru/Waterloo-Hawaii-IfA G-band Survey (WHIGS, g), The Panoramic Survey Telescope & Rapid Response System (Pan-STARRS, i), and Subaru/Wide Imaging with Subaru HSC of the Euclid Sky (WISHES, z). The main UNIONS footprint is located at a declination (DEC.) $30^{\circ}$ and a galactic latitude of $|b|>25^{\circ}$, enclosing a total area of $4,861\mathrm{\ deg^{2}}$. Extended $u$- and $z$- (WISHES+) observations are being conducted to cover the sky area at $\mathrm{DEC.}<30^{\circ}$, elevating the sky coverage of UNIONS. Albeit UNIONS is motivated by the Euclid wide survey (euclid_collaboration_euclid_2022), efforts have been devoted to maintaining UNIONS as an independent survey with scientific values beyond what will be brought about by the Euclid.
@@
The $r$-band observation of CFHT/CFIS has an excellent median seeing of $0.69\arcsec$ and reaches a surface brightness limit of 28.4 $\mathrm{mag/arcsec^{2}}$ for extended sources. It is set as the reference catalog by which the elliptical aperture is defined for the forced photometry applied to all five UNIONS bands. Since the broadband observations of UNIONS are conducted under multiple telescopes, the Gaussian Aperture and Photometry ($\mathtt{GAaP}$; kuijken_gravitational_2015) is adopted to produce Gaussianized images for raw imaging data collected from observations at different instruments. The forced photometry with Gaussian-weighted, elliptical apertures can then be properly used to perform homogeneous multiband photometric measurements simultaneously. The depth of UNIONS is characterized by the $10\sigma$ point source depth measured through a $2.0\arcsec$ aperture. The current median depth of UNIONS is $u,g,r,i,{z}_{\mathrm{PS}},{z}_{\mathrm{HSC}}]=[23.7,24.5,24.2,23.8,23.3,23.3]$, to be compared with the original goal of $[u,g,r,i,z]=[23.6,24.5,23.9,23.6,23.4]$ (gwyn_unions_2025).
In this work, we use the latest internal 5-band data release UNIONS5000 (as of May 2025; $\mathrm{UNIONS5000}_{ugriz}$ hereafter) that includes $\sim 20,600$ tiles with a total processed survey area of $\sim 4,200\mathrm{\ deg^{2}}$, covering both north (NGC) and south Galactic caps (SGC). $\mathrm{UNIONS5000}_{ugriz}$ contains photometric measurements using forced photometry adopting an aperture size of $\mathrm{APER}=0.7\arcsec$ and $\mathrm{APER}=1.0\arcsec$, and an additional one with the optimal aperture $\mathrm{APER_{opt}}$. In most cases, the optimal one is equal to the smallest possible aperture of $\mathrm{APER}=0.7\arcsec$. In the case where the measurement adopting $\mathrm{APER}=0.7\arcsec$ fails, the optimal photometric measurement is then that using $\mathrm{APER}=1.0\arcsec$ (see also kuijken_fourth_2019). We always use the optimal photometry for the data presented in this work.
Figure 2: The angular separation-$\mathtt{NN\_dist}$ plane for UNIONS-VLASS sources cross-matched with a search radius of $6^{\prime\prime}$. The vertical lines indicate the $\mathtt{NN\_dist}$ cut of $30^{\prime\prime}$ and the horizontal line represents search radius of $1.5^{\prime\prime}$ to perform the cross-match (see § 3.2 and 3.3).
2.2 VLASS